40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3094     Entry time: Mon Jun 21 18:01:34 2010
Author: Jenne 
Type: Update 
Category: 40m Upgrading 
Subject: SRM, PRM hung, magnets inspected 

[Jenne, Kiwamu, Rana, Eric Gustafson]

The SRM and PRM have been re-hung, and are ready for installation into the chambers.  Once we put the OSEMs in, we may have to check the rotation about the Z-axis.  That was not confirmed today (which we could do with the microscope on micrometer, or by checking the centering of the magnets in the OSEMs).

Also, Eric and Rana inspected the Tip Tilt magnets, and took a few that they did their best to destroy, and they weren't able to chip the magnets.  There was concern that several of the magnets showed up with the coatings chipped all over the place.  However, since Rana and Eric did their worst, and didn't put any new chips in, we'll just use the ones that don't have chips in them.  Rana confiscated all the ones with obvious bad chips, so we'll check the strengths of the other magnets using a gaussmeter, and choose sets of 4 that are well matched. 

Eric, photographer extraordinaire, will send along the pictures he took, and we'll post them to Picasa.

ELOG V3.1.3-